OpenStudy (anonymous):

What 2 integers that make the equation y^2=25? 5, –5 5, 0.5 5, 25 i will post the last answer in a sec

OpenStudy (anonymous):

first two

OpenStudy (anonymous):

\[(-5)^2=25\] and \[5^2=25\]

OpenStudy (anonymous):

\[-\sqrt{5}, \sqrt{5}\] the last answer

OpenStudy (anonymous):

k

OpenStudy (anonymous):

thanks!